# Enhancing Real-World Complex Network Representations with Hyperedge Augmentation

### v1

@@ 3.1. Hyperedge Construction
Figure 2. Graph statistics-based HyperAug can help GNNs surpass the limitations posed by the 1-WL test. (a) Without HyperAug, GNNs are unable to distinguish the two example non-isomorphic graphs that have the same node degrees; (b) With HyperAug, GNNs can now successfully distinguish those graphs, since HyperAug identifies that the right graph contains two cliques, while the left graph contains none.
From a different Data Perspective.
@@
Many graphs carry information from different data perspectives. Stacking various types of information from different data perspectives, whether as the graph’s node features or edges, can potentially make the learning challenging, since the underlying GNN would then have to learn from these manually mixed information. Real-world examples of such graphs include the biological networks, where both entity-entity interactions and geometrical information are critical to the prediction tasks. For example, gene regulatory interactions might be represented as an adjacency matrix, while the spatial positioning of genes can be used to establish higher-order relationships. The data inherently offer multiple ways for connecting entities within a graph. In this scenario, HyperAug constructs virtual hyperedges by extracting information, preferably featuring multi-node relations, from a different data perspective than the node embeddings and edges. These hyperedges can then be processed separately from the node embeddings and edge adjacency, which are derived from a different data perspective, and ultimately serve as auxiliary node embeddings for the prediction tasks. This type of virtual hyperedge augmentation serves not only as a complement to to the simple edges, but also as an enhancement to the overall information obtained – this additional information is from a different data source that contain high-order node realtion that cannot be simply captured by the simple edges or node embeddings of the original graph.
From a different Data Modality.
@@
For graphs that incorporate information from various data modalities, such as product networks containing both textual and image data, effectively integrating the information from different data modalities can be challenging for a GNN. Similar to the previous scenario, HyperAug addresses this by constructing virtual hyperedges using information from one of the data modalities, preferably exhibiting multi-node relations, which are distinct from the other data modalities used by node embeddings and edges. With this approach, information from different data modalities can be processed independently, and the data from the additional modality essentially serve as an augmentation. Note that this differs from simply incorporating an additional data source into the node embedding, as the alternative data modality may only convey ‘grouping’ information, which is more suitably represented as higher-order graph relations, such as virtual hyperedges. As the virtual hyperedges are constructed from a different modality, this type of virtual hyperedge augmentation provides both a complement to the simple edges of the original graph, and an enhancement to the overall information.
@@
In graph representation learning, while the simple edges of the graph can be represented as an adjacency matrix $\mathbf{A}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $A_{uv}=1$ if $(u,v)\in\mathcal{E}$ and 0 otherwise, the virtual hyperedges of the augmented graph can be represented as an incidence matrix $\mathbf{H}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}_{h}|}$, with each entry $H_{ve_{h}}=1$ if $v\in e_{h}$ and 0 otherwise. The augmented graph $\mathcal{G}^{\prime}$ is then further processed to produce the final auxiliary features for the original graph’s nodes.
3.2. Utilising the Hyperedge Information
In order to thoroughly and appropriately utilise the additional information from HyperAug’s virtual hyperedges, we introduce a feature extraction function $f$ that operates specifically on the augmented graph with virtual hyperedges $\mathcal{G}^{\prime}$, and outputs a set of auxiliary node features $\mathbf{Z}^{\prime}$:
(1) $\mathbf{Z}^{\prime}=f(\mathcal{G}^{\prime},\mathbf{X},\mathbf{E})$
where $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d_{v}}$ denotes the node feature matrix of the graph, with each row $\mathbf{x}_{v}\in\mathbb{R}^{d_{v}}$ being the $d_{v}$-dimensional features of node $v$, and $\mathbf{E}\in\mathbb{R}^{|\mathcal{E}|\times d_{e}}$ denotes the edge feature matrix of the graph, with each row $\mathbf{e}_{e}\in\mathbb{R}^{d_{e}}$ being the $d_{e}$-dimensional features of edge $e$. To integrate the auxiliary features $\mathbf{Z}^{\prime}$ into the node embeddings $\mathbf{Z}$ obtained from the original graph $\mathcal{G}$, we employ a simple mechanism: concatenating the original node embeddings $\mathbf{Z}$ and the auxiliary features $\mathbf{Z}^{\prime}$, regularised by a linear layer with weights $\bm{\theta}$ and bias $\mathbf{b}$, as defined in the following equation:
(2) $\mathbf{Z}_{\text{aug}}=\bm{\theta}\cdot(\mathbf{Z}\>\|\>\mathbf{Z}^{\prime})+\mathbf{b}$
The final augmented node embeddings $\mathbf{Z}_{\text{aug}}$ are then used for the downstream prediction tasks. The complete HyperAug pipeline is illustrated in Figure 1. While this flexible HyperAug pipeline allows other mechanisms for integrating the auxiliary features into the original graph than using a linear layer, it is worth pointing out that this work primarily focuses on exploring the potential of leveraging GNN performance through incorporating higher-order node relations into the original graphs. Designing fine-grained feature integration mechanisms is not the main focus of this research.
4. Experiments
4.1. Datasets
@@
Table 2. Accuracy (%) of the HyperAug variants, whose auxiliary features are integreted at different stages, compared with the vanilla GCN without any augmentation. The GCN with HyperAug($f$, emb.) variants show a consistent and much more significant accuracy enhancement compared with the vanilla GCN and the GCN with HyperAug($f$, input) variants.
@@
GCN 87.2 $\pm$ 0.0 62.5 $\pm$ 0.0 75.6 $\pm$ 4.1
GCN+HyperAug(HyperConv, input) 86.9 $\pm$ 0.3 63.6 $\pm$ 1.3 96.0 $\pm$ 0.5
GCN+HyperAug(ED-HNN, input) 87.1 $\pm$ 0.5 63.6 $\pm$ 1.3 95.8 $\pm$ 0.4
GCN+HyperAug(HyperConv, emb.) 87.2 $\pm$ 0.0 63.7 $\pm$ 0.2 96.8 $\pm$ 0.5
GCN+HyperAug(ED-HNN, emb.) 87.4 $\pm$ 0.3 66.7 $\pm$ 2.6 98.1 $\pm$ 0.7
To our knowledge, there is a deficiency in the quantity and variety of widely-accepted graph datasets that support data augmentation beyond simple graphs. This is largely attributable to the scarcity of explicitly labelled higher-order node relationships. In order to thoroughly evaluate the effectiveness of HyperAug across varied domains, we build 23 novel graph datasets derived from real-world networks across varied domains, including social media, biology, and e-commerce. Care has been taken to ensure that the datasets do not contain any personally identifiable information. The 23 datasets can be splitted into three groups according to the intended virtual hyperedge construction process for HyperAug:
MUSAE
We build eight social networks derived from the Facebook pages, GitHub developers and Twitch gamers, plus three English Wikipedia page-page networks on specific topics (chameleons, crocodiles and squirrels) based on MUSAE (Rozemberczki et al., 2021). Nodes represent users or articles, and edges are mutual followers relationships between the users, or mutual links between the articles. These datasets are intended to assess HyperAug’s effectiveness in constructing virtual hyperedges from the graph statistics: on these datasets, HyperAug constructs the virtual hyperedges to be mutually connected sub-groups that contain at least three nodes (i.e., maximal cliques with sizes of at least 3).
GRAND
We select and build ten gene regulatory networks in different tissues and diseases from GRAND (Ben Guebila et al., 2022), a public database for gene regulation. Nodes represent gene regulatory elements (Maston et al., 2006), and edges are regulatory effects between genes. Node features are extracted based on the gene sequences. These datasets are intended to assess HyperAug’s effectiveness in constructing virtual hyperedges from a different data perspective: on these datasets, HyperAug constructs the virtual hyperedges by grouping geometrically nearby genomic elements on the chromosomes, i.e., the genomic elements within 200k base pair distance.
Amazon
Following existing works on graph representation learning on e-commerce networks (Shchur et al., 2018; Zeng et al., 2020), we faithfully reconstruct a subset of the OGB (Hu et al., 2020) ogbn-products dataset, and build two product co-purchase/co-review networks based on the Amazon Product Reviews dataset (McAuley et al., 2015; He and McAuley, 2016; Ni et al., 2019). Nodes represent products, and an edge between two products is established if a user buys or writes reviews for both products. Node features are extracted based on the textual description of the products. These datasets are intended to assess HyperAug’s effectiveness in constructing virtual hyperedges from a different data modality: on these datasets, HyperAug introduces the image modality into the construction of virtual hyperedges. To be specific, the raw images of the products are fed into a CLIP (Radford et al., 2021) classifier, and a 512-dimensional feature embedding for each image is returned to assist the clustering. HyperAug then constructs the virtual hyperedges by grouping products whose image embeddings have pairwise distances within a certain threshold.
In addition to the 23 novel datasets, we also adopt three commonly used citation datasets: Cora-CoCitation, Cora-CoAuthorship, and Pubmed-CoCitation, to match HyperAug’s performance with the community standard. We use these three datasets to also assess HyperAug’s effectiveness in constructing virtual hyperedges from a different data perspective: the edges of the original graph and the virtual hyperedges constructed by HyperAug are co-citation links and co-authorship groups respectively, or vice versa (i.e., co-authorship links and co-citation groups). Table 1 reports the key graph statistics for each dataset group, and more details of the used datasets are described in Appendix B. We make our source code and full datasets publicly available at https://anonymous-url/.
4.2. Designing HyperAug: Placement Choices and Feature Extraction Techniques
Table 3. Accuracy (%) of HyperAug with various simple-GNNs and hyper-GNNs as feature extraction functions, compared with the vanilla GCN without any augmentation. The HyperAug variants with hyper-GNNs as feature extraction functions clearly outperform HyperAug with simple-GNNs.
@@ 4.2. Designing HyperAug: Placement Choices and Feature Extraction Techniques
GCN 87.2 $\pm$ 0.0 62.5 $\pm$ 0.0 75.6 $\pm$ 4.1
GCN+HyperAug(GAT) 86.7 $\pm$ 0.1 62.5 $\pm$ 0.3 91.3 $\pm$ 0.1
GCN+HyperAug(GraphSAGE) 87.0 $\pm$ 0.4 65.3 $\pm$ 0.9 93.0 $\pm$ 0.0
GCN+HyperAug(HyperConv) 87.2 $\pm$ 0.0 63.7 $\pm$ 0.2 96.8 $\pm$ 0.5
GCN+HyperAug(ED-HNN) 87.4 $\pm$ 0.3 66.7 $\pm$ 2.6 98.1 $\pm$ 0.7
4.2.1. Appropriate Phase for Applying HyperAug
Before conducting the full evaluation on HyperAug, we analyse the optimal usage of HyperAug through a series of ablation studies. The auxiliary features generated by HyperAug provide significant flexibility regarding the stage of integration, including simultaneous injection with the input or blending with the original GNN’s activations. To thoroughly identify the appropriate place for inserting HyperAug’s auxiliary features, we compare the performance of HyperAug with the following settings: (1) concatenate HyperAug’s auxiliary features directly with the input, proceed the concatenated features with a linear layer, then feed into the GNN (denoted as HyperAug($f$, input), where $f$ is the feature extraction function); (2) concatenate HyperAug’s auxiliary features with the GNN activations, followed by a linear layer (denoted as HyperAug($f$, emb.)). This ablation study is conducted on the MUSAE-GitHub, GRAND-Brain and Amazon-Computers datasets, each featuring a distinct virtual hyperedge construction strategy. These augmetnations are then used on top of a GCN (Kipf and Welling, 2017). HyperConv (Bai et al., 2021) and ED-HNN (Wang et al., 2023) are selected as the feature extraction functions for HyperAug, since HyperConv is one of the most classic hyper-GNN baseline, and ED-HNN is known as the state-of-the-art in hyper-GNNs up to the point of this research.
Table 2 summarises the results for this set of ablation studies. The results show a consistent and much more significant increase in accuracy for GCN with HyperAug($f$, emb.), compared with the vanilla GCN and GCN with HyperAug($f$, input), demonstrating that inserting HyperAug’s auxiliary features at the output embedding phase is the more appropriate choice. Inserting HyperAug’s auxiliary features at the input phase does not guarantee to leverage GCN’s performance. This is because concatenating the auxiliary features with the node features at the input phase can potentially mix-up those different types of information, and confuses the GNNs. This also contradicts the intuition behind HyperAug.
4.2.2. Choice of the Feature Extraction Function in HyperAug
It is intuitive to use hyper-GNNs as feature extraction functions in HyperAug, so that the virtual hyperedge information can be fully captured. However, it is still worth verifying that hyper-GNNs indeed outperform simple-GNNs as feature extraction functions in HyperAug, in order to prove the efficacy of HyperAug. Therefore, we compare the performance of HyperAug with different classes of GNNs as the feature extraction function $f$, denoted as HyperAug($f$): (1) hyper-GNNs, namely HyperConv and ED-HNN; (2) simple-GNNs, namely GAT (Veličković et al., 2018) and GraphSAGE (Hamilton et al., 2017). These ablation experiments are again conducted on the MUSAE-GitHub, GRAND-Brain and Amazon-Computers datasets, and GCN remains as the GNN for applying HyperAug. Starting from this stage, all experiments apply HyperAug’s auxiliary features at the output embedding phase. Table 3 reports the results for this set of ablation studies. The results clearly show that HyperAug with hyper-GNNs as feature extraction functions outperform HyperAug with simple-GNNs, thereby justifying that the HyperAug pipeline indeed utilises the virtual hyperedges it constructs.
Table 4. Average accuracy rankings of three vanilla GNNs and HyperAug with various feature extraction functions on different GNNs, across 23 node classification datasets. GraphSAGE with HyperAug using HyperConv as the feature extraction function proves to be the best-performing model.

### v2

@@ 3.1 Combinatorial Complex Construction
Figure 2: Graph statistics-based TopoAug can help GNNs surpass the limitations posed by the 1-WL test. (a) Without TopoAug, GNNs are unable to distinguish the two example non-isomorphic graphs that have the same node degrees; (b) Since TopoAug identifies that the right graph contains two cliques, while the left graph contains none, GNNs can now successfully distinguish those graphs.
From the Graph Statistics
For graphs that can reveal valuable information from their local clusters and cliques, such as the social networks, TopoAug constructs hyperedges by computing from graph statistics, such as maximal cliques. TopoAug then groups those nodes within the same maximal cliques. This augmentation acts as a complement to the simple edges in the original graph by calculating the local node-wise clustering relationships. TopoAug then enables GNNs seeking to utilise local clustering information to rapidly access that information without needing to explicitly precompute it. In addition, by separating out the local clusters and cliques from the simple edge adjacency of the original graph, TopoAug can prevent GNNs from mixing up those information, thereby potentially enhancing its training performance. Another crucial advantage for this type of augmentation is that it enables GNNs to overcome the expressiveness limitations posed by the 1-WL test [14, 37, 38]: it is able to help GNNs distinguish non-isomorphic graphs with identical node degrees, by specifying their distinct maximal cliques, as illustrated in Figure 2.
From a different Data Perspective
@@
Many graphs carry information from different data perspectives. Stacking various types of information from different data perspectives, whether as the graph’s node features or edges, can potentially make the learning challenging, since the underlying GNN would then have to learn from these manually mixed information. Real-world examples of such graphs include the biological networks, where both entity-entity interactions and geometrical information are critical to the prediction tasks. For example, gene regulatory interactions might be represented as an adjacency matrix, while the spatial positioning of genes can be used to establish higher-order relationships. The data inherently offer multiple ways for connecting entities within a graph. In this scenario, TopoAug constructs hyperedges by extracting information, preferably featuring multi-node relations, from a different data perspective than the node embeddings and edges. These hyperedges can then be processed separately from the node embeddings and edge adjacency, which are derived from a different data perspective, making them ultimately serve as auxiliary node embeddings for the prediction tasks. This type of virtual hyperedge augmentation serves not only as a complement to the simple edges, but also as an enhancement to the overall information obtained – it takes additional information from a different data source that contain high-order node relations.
From a different Data Modality
@@
For graphs that incorporate information from various data modalities, such as product networks containing both textual and image data, effectively integrating the information from different data modalities can be challenging for a GNN. Similar to the previous scenario, TopoAug addresses this by constructing hyperedges using information from one of the data modalities, preferably exhibiting multi-node relations, which are distinct from the other data modalities used by node embeddings and edges. With this approach, information from different data modalities can be processed independently, and the data from the additional modality essentially serve as an augmentation. Note that this differs from simply incorporating an additional data source into the node embedding, as the alternative data modality may only convey ‘grouping’ information, which is more suitably represented as higher-order graph relations, such as hyperedges. As the hyperedges are constructed from a different modality, this type of hyperedge augmentation provides both a complement to the simple edges of the original graph, and an enhancement to the overall information.
@@
In topological deep learning, while the simple edges of the graph can be represented as an adjacency matrix $\mathbf{A}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $A_{uv}=1$ if $(u,v)\in\mathcal{E}$ and 0 otherwise, the hyperedges, as well as the final augmented combinatorial complex, can be represented as incidence matrices $\mathbf{H}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}_{h}|}$ and $\mathbf{H}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{X}|}$ respectively, with each entry $H_{ve_{h}}\>(H_{vx})=1$ if $v\in e_{h}\>(x)$ and 0 otherwise. The combinatorial complex $\mathcal{X}$ is then further processed to produce the final auxiliary features for the original graph’s nodes.
3.2 Utilising the Hyperedge Information
In order to thoroughly and appropriately utilise the additional information from TopoAug’s combinatorial complex, we introduce a node feature augmentation pipeline consisting of:
  1. 1.
    A combinatorial complex construction function $f_{\text{CC}}:\mathfrak{G}\to\mathfrak{X}$ mapping from the graph space $\mathfrak{G}$ to the combinatorial complex space $\mathfrak{X}$. $f_{\text{CC}}$ also encapsulates a hyperedge construction function $h$ (defined in Section 3.1); followed by
  2. 2.
    A node embedding function $f_{\text{emb}}:\mathfrak{G}\times\mathbb{R}^{|\mathcal{V}|\times d_{v}}\times\mathbb{R}^{|\mathcal{E}|\times d_{e}}\to\mathbb{R}^{|\mathcal{V}|\times d_{z}}$ that operates on the original graph, together with its unaugmented node and edge features, to compute the $d_{z}$-dimensional original node embeddings; and
  3. 3.
    An auxiliary function $f_{\text{aux}}:\mathfrak{X}\times\mathbb{R}^{|\mathcal{V}|\times d_{v}}\times\mathbb{R}^{|\mathcal{E}|\times d_{e}}\to\mathbb{R}^{|\mathcal{V}|\times d_{z^{\prime}}}$ that operates on the constructed combinatorial complex, together with the original graph’s unaugmented node and edge features to produce the $d_{z^{\prime}}$-dimensional auxiliary node features.
The pipeline outputs the final, $d_{z_{\text{aug}}}$-dimensional augmented node embeddings as defined in the following equations:
$\text{TopoAug}(\mathcal{G},\mathbf{X},\mathbf{E})=(f_{\text{emb}}(\mathcal{G},\mathbf{X},\mathbf{E})\>\|\>f_{\text{aux}}(f_{\text{CC}}(\mathcal{G}),\mathbf{X},\mathbf{E}))\mathbf{W}$ (5)
[…]
4 Experiments
4.1 Datasets
Table 1: Aggregated statistics of the datasets used for TopoAug evaluation.
Name Hyperedge Construction Mechanism #Datasets Avg. #Nodes Avg. #Edges Avg. #Hyperedges Avg. Node Degree Avg. Hyperedge Degree #Classes
MUSAE-GitHub Graph Statistics 1 37,700 578,006 223,672 30.7 4.6 4
MUSAE-Facebook Graph Statistics 1 22,470 342,004 236,663 30.4 9.9 4
MUSAE-Twitch Graph Statistics 6 5,686 143,038 110,142 50.6 6.0 2
MUSAE-Wiki Graph Statistics 3 6,370 266,998 118,920 88.8 14.4 Regression
GRAND-Tissues Multi-Perspective 6 5,931 5,926 11,472 2.0 1.3 3
GRAND-Diseases Multi-Perspective 4 4,596 6,252 7,743 2.7 1.3 3
Cora Multi-Perspective 2 2,708 5,429 1,326 4.0 3.5 7
Pubmed Multi-Perspective 1 19,717 44,338 7,963 4.5 4.3 3
Amazon-Computers Multi-Modality 1 10,226 55,324 10,226 10.8 4.0 10
Amazon-Photos Multi-Modality 1 6,777 45,306 6,777 13.4 4.8 10
To the best of our knowledge, there is a deficiency in the quantity and variety of widely-accepted graph datasets that support data augmentation beyond simple graphs. This is largely attributable to the scarcity of explicitly labelled higher-order node relationships. In order to thoroughly evaluate the effectiveness of TopoAug across varied domains, we build 23 novel graph datasets derived from real-world networks across varied domains, including social media, biology, and e-commerce. Care has been taken to ensure that the datasets do not contain any personally identifiable information. The 23 datasets can be split into three groups according to the intended virtual hyperedge construction process for TopoAug:
MUSAE
We build eight social networks derived from the Facebook pages, GitHub developers and Twitch gamers, plus three English Wikipedia page-page networks on specific topics (chameleons, crocodiles and squirrels) based on MUSAE [29]. Nodes represent users or articles, and edges are mutual followers relationships between the users, or mutual links between the articles. These datasets are intended to assess TopoAug’s effectiveness in constructing virtual hyperedges from the graph statistics: on these datasets, TopoAug constructs the virtual hyperedges to be mutually connected sub-groups that contain at least three nodes (i.e., maximal cliques with sizes of at least 3). The tasks for the Facebook, GitHub, and Twitch datasets involve multi-class classification to predict the categories of users or pages, while the task for the Wiki dataset is a regression task that predicts the average monthly traffic of a web page.
GRAND
We select and build ten gene regulatory networks in different tissues and diseases from GRAND [2], a public database for gene regulation. Nodes represent gene regulatory elements [23] with three distinct types: protein-encoding gene, lncRNA gene [22], and other regulatory elements. Edges are regulatory effects between genes. We train a CNN [8] and use it to take the gene sequence as input and create a 4,651-dimensional embedding for each node. These datasets are intended to assess TopoAug’s effectiveness in constructing virtual hyperedges from a different data perspective: on these datasets, TopoAug constructs the virtual hyperedges by grouping geometrically nearby genomic elements on the chromosomes, i.e., the genomic elements within 200k base pair distance. The task is a multi-class classification of gene regulatory elements.
Amazon
Following existing works on graph representation learning on e-commerce networks [31, 41], we faithfully reconstruct a subset of the OGB [17] ogbn-products dataset, and build two product co-purchase/co-review networks based on the Amazon Product Reviews dataset [16, 24, 26]. Nodes represent products, and an edge between two products is established if a user buys or writes reviews for both products. Node features are extracted based on the textual description of the products. These datasets are intended to assess TopoAug’s effectiveness in constructing virtual hyperedges from a different data modality: on these datasets, TopoAug introduces the image modality into the construction of virtual hyperedges. To be specific, the raw images of the products are fed into a CLIP [27] classifier, and a 512-dimensional feature embedding for each image is returned to assist the clustering. TopoAug then constructs the virtual hyperedges by grouping products whose image embeddings have pairwise distances within a certain threshold. The task is to predict the sub-category of a product in a multi-class classification setup.
In addition to the 23 novel datasets, we also adopt three commonly used citation datasets: Cora-CoCitation, Cora-CoAuthorship, and Pubmed-CoCitation, to match TopoAug’s performance with the community standard. We use these three datasets to also assess TopoAug’s effectiveness in constructing virtual hyperedges from a different data perspective: the edges of the original graph and the virtual hyperedges constructed by TopoAug are co-citation links and co-authorship groups respectively, or vice versa (i.e., co-authorship links and co-citation groups). Table 1 reports the key graph statistics for each dataset group, and more details of the datasets are described in Appendices A and B. We make our source code and full datasets publicly available at https://github.com/VictorZXY/TopoAug.
4.2 Experimental Setup
Training Details
We run all the experiments on NVIDIA A100 and V100 GPUs, with up to 40GB memory. Adam [19] is used as the optimiser, and CosineAnnealingLR [11] is used as the learning rate scheduler for all training. For each experiment, the nodes of the graph dataset are randomly split into training, validation, and test sets with a split ratio of 6:2:2. All models are trained for 500 epochs. For node classification tasks, the negative log likelihood loss (NLLLoss) is used as the loss function. For node regression tasks, the mean square error (MSELoss) is used as the loss function. Each experiment typically takes less than 5 minutes to train, when ED-HNN is not incorporated in the model. Due to the significantly larger architecture of ED-HNN, experiments involving it can take up to 2 hours.
Hyperparameter Settings
We perform a hyperparameter search for the learning rate and dropout rate, while keeping the hidden dimension of the layers fixed as 64. To ensure fair comparison, all evaluated GNNs (GCN, GAT, GraphSAGE, HyperConv, and ED-HNN) share the same hyperparameter combinations. After hyperparameter searching, we adopt the following hyperparameter selections: learning rate = 0.001, and dropout rate = 0.5. For the additional hyperparameters of ED-HNN, we closely adhere to the hyperparameter settings specified in the original ED-HNN paper [34], and set the number of all inner multi-layer perceptrons (MLPs) within ED-HNN to 2.
4.3 Designing TopoAug
As described in Section 3.2, the TopoAug pipeline consists of a combinatorial complex construction function $f_{\text{CC}}$ that constructs a combinatorial complex from the original graph, followed by a backbone node embedding function $f_{\text{emb}}$ that operates on the original graph to produce the unaugmented node embeddings, and an auxiliary function $f_{\text{emb}}$ that operates on the constructed combinatorial complex to produce the auxiliary node features. This opens up the following questions regarding the practical design of TopoAug:
  • At what point should the auxiliary node features be integrated with the original node information – directly at the input side, or within the middle-layer embeddings?
  • What is the appropriate type, and consequently, the optimal choice of the auxiliary function?
We investigate these design choices and analyse the optimal usage of TopoAug through a series of ablation studies.
4.3.1 Appropriate Phase for Applying TopoAug
The auxiliary features generated by TopoAug provide significant flexibility regarding the stage of integration, including simultaneous injection with the input node features or blending with the original GNN’s activations. To thoroughly identify the appropriate place for inserting TopoAug’s auxiliary features, we compare the performance of TopoAug with the following settings:
  1. 1.
    concatenate TopoAug’s auxiliary features directly with the input node features, then proceed the concatenated features with a linear layer, before feeding into the embedding GNN (denoted as TopoAug($f_{\text{aux}}$, input), where $f_{\text{aux}}$ is the auxiliary function);
  2. 2.
    concatenate TopoAug’s auxiliary features with the GNN activations, followed by a linear layer (denoted as TopoAug($f_{\text{aux}}$, emb.)).
We then conducted experiments on the MUSAE-GitHub, GRAND-Brain and Amazon-Computers datasets, each featuring a distinct virtual hyperedge construction strategy, to evaluate these two settings. The augmentations are applied to a GCN [20]. HyperConv [1] and ED-HNN [34] are selected as the auxiliary functions for TopoAug, since HyperConv is one of the most popular hyper-GNN baselines, and ED-HNN represents the current state-of-the-art in hyper-GNNs.
Table 2 summarises the results for this set of ablation studies. The results show a consistent and significant increase in accuracy for GCN with TopoAug($f_{\text{aux}}$, emb.) compared to the vanilla GCN and also GCN with TopoAug($f_{\text{aux}}$, input). This result suggests that inserting TopoAug’s auxiliary features at the output embedding phase is the more appropriate choice. Intuitively, this is because concatenating the auxiliary features with the node features at the input phase can potentially mix up different types of input information, and confuse the GNN.
@@ 4.3.1 Appropriate Phase for Applying TopoAug
Table 2: Evaluating different phases for applying TopoAug (TopoAug($f_{\text{aux}}$, input), TopoAug($f_{\text{aux}}$, emb.)): accuracy (%) of the TopoAug variants using GCN as the backbone embedding function $f_{\text{emb}}$, whose auxiliary features are integrated at different stages, compared with the vanilla GCN without any augmentation. GCN with TopoAug($f_{\text{aux}}$, emb.) variants show a consistent and much more significant accuracy enhancement compared with vanilla GCN and GCN with TopoAug($f_{\text{aux}}$, input) variants.
@@
GCN baseline 87.2 $\pm$ 0.0 62.5 $\pm$ 0.0 75.6 $\pm$ 4.1
TopoAug(HyperConv, input) 86.9 $\pm$ 0.3 63.6 $\pm$ 1.3 96.0 $\pm$ 0.5
TopoAug(ED-HNN, input) 87.1 $\pm$ 0.5 63.6 $\pm$ 1.3 95.8 $\pm$ 0.4
TopoAug(HyperConv, emb.) 87.2 $\pm$ 0.0 63.7 $\pm$ 0.2 96.8 $\pm$ 0.5
TopoAug(ED-HNN, emb.) 87.4 $\pm$ 0.3 66.7 $\pm$ 2.6 98.1 $\pm$ 0.7
Table 3: Evaluating different types for the auxiliary function $f_{\text{aux}}$: accuracy (%) of TopoAug with various simple-GNNs and hyper-GNNs as auxiliary functions, using GCN as the backbone embedding function $f_{\text{emb}}$, compared with the vanilla GCN without any augmentation. The TopoAug variants with hyper-GNNs as auxiliary functions clearly outperform TopoAug with simple-GNNs.
@@
GCN baseline 87.2 $\pm$ 0.0 62.5 $\pm$ 0.0 75.6 $\pm$ 4.1
TopoAug(GAT) 86.7 $\pm$ 0.1 62.5 $\pm$ 0.3 91.3 $\pm$ 0.1
TopoAug(GraphSAGE) 87.0 $\pm$ 0.4 65.3 $\pm$ 0.9 93.0 $\pm$ 0.0
TopoAug(HyperConv) 87.2 $\pm$ 0.0 63.7 $\pm$ 0.2 96.8 $\pm$ 0.5
TopoAug(ED-HNN) 87.4 $\pm$ 0.3 66.7 $\pm$ 2.6 98.1 $\pm$ 0.7
4.3.2 Choice of the Auxiliary Function Type
It is intuitive to use hyper-GNNs as auxiliary functions ($f_{\text{aux}}$) in TopoAug, so that hyperedge information can be fully captured. However, it is still worth verifying that hyper-GNNs indeed outperform simple-GNNs as auxiliary functions in TopoAug, in order to prove the efficacy of TopoAug. Therefore, we compare the performance of TopoAug with different classes of GNNs as the auxiliary function, denoted as TopoAug($f_{\text{aux}}$): (1) hyper-GNNs, namely HyperConv and ED-HNN; (2) simple-GNNs, namely GAT [33] and GraphSAGE [13]. These ablation experiments are again conducted on the MUSAE-GitHub, GRAND-Brain and Amazon-Computers datasets, and GCN remains as the GNN for applying TopoAug. Starting from this stage, all experiments apply TopoAug’s auxiliary features at the output embedding phase. Table 3 reports the results for this set of ablation studies, which clearly show that TopoAug with hyper-GNNs as auxiliary functions outperform TopoAug with simple-GNNs, thereby justifying that TopoAug indeed utilises the virtual hyperedges it constructs.
4.3.3 Identifying the Best Auxiliary Function
Table 4: Evaluating the optimal auxiliary function $f_{\text{aux}}$ for TopoAug: average accuracy rankings of three vanilla GNNs and TopoAug with various auxiliary functions on different GNNs, across 23 node classification datasets. GraphSAGE with TopoAug using HyperConv as the auxiliary function proves to be the best-performing model.